### v1

# Large Language Model Can Continue Evolving From Mistakes

Abstract
Large Language Models (LLMs) demonstrate impressive performance in various downstream tasks. However, they may still generate incorrect responses in certain scenarios due to the knowledge deficiencies and the flawed pre-training data. Continual Learning (CL) is a commonly used method to address this issue. Traditional CL is task-oriented, using novel or factually accurate data to retrain LLMs from scratch. However, this method requires more task-related training data and incurs expensive training costs. To address this challenge, we propose the Continue Evolving from Mistakes (CEM) method, inspired by the ’summarize mistakes’ learning skill, to achieve iterative refinement of LLMs. Specifically, the incorrect responses of LLMs indicate knowledge deficiencies related to the questions. Therefore, we collect corpora with these knowledge from multiple data sources and follow it up with iterative supplementary training for continuous, targeted knowledge updating and supplementation. Meanwhile, we developed two strategies to construct supplementary training sets to enhance the LLM’s understanding of the corpus and prevent catastrophic forgetting. We conducted extensive experiments to validate the effectiveness of this CL method. In the best case, our method resulted in a 17.00% improvement in the accuracy of the LLM.
Large Language Model Can Continue Evolving From Mistakes
Haokun Zhao, Haixia Han, Jie Shi, Chengyu Du, Jiaqing Liang, Yanghua Xiao Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University hkzhao23@m.fudan.edu.cn
1 Introduction
Figure 1: Continual evaluation can help us uncover knowledge deficiency of the LLM and thus evolve the model through continuous learning.
With the exponential expansion of both data and parameter size, large language models (LLMs) demonstrate outstanding performance across a series of downstream tasks Chang et al. (2023). However, LLMs have the potential to generate incorrect outputs, a phenomenon often referred to as ’hallucination’ Huang et al. (2023). In addition, it is important to note that satisfactory answers generated by LLMs cannot be achieved within a single training session Han et al. (2024). Continual refinement and optimization processes are necessary. This optimization process ensures that the LLM can undergo iterative refinement, continuously enhancing its performance.
There are two main reasons for the poor performance of LLM on specific questions. Firstly, some of these inquiries exceed the knowledge boundaries of LLM, indicating either a deficiency in domain expertise or outdated information within the LLM Ren et al. (2023); Zhang et al. (2023). Secondly, the LLM may incorporate erroneous knowledge derived from potentially unreliable sources in the training data, resulting in the generation of inaccuracies or misconceptions that contradict factual information Rae et al. (2021); Zhao et al. (2023).
To rectify the model’s knowledge deficits and ensure alignment with contemporary world knowledge, a common method is periodic training from scratch using the domain or latest corpus, such as the most recently archived Wikipedia. However, this method incurs significant training expenses Patterson et al. (2021). Another method involves pre-training on a corpus that includes accurate information. However, this method may result in a decline in other aspects of the model’s performance Li et al. (2022). This challenge has prompted research into mitigating catastrophic forgetting in continuous learning. Existing studies can be roughly grouped into three categories, e.g., experience replay methods Chaudhry et al. (2019); Wu et al. (2021), regularization-based methods Kirkpatrick et al. (2017a); Chen et al. (2022), and dynamic architecture methods Mallya et al. (2018). Recently, researchers have designed some hybrid methods that take advantage of the aforementioned techniques Chen et al. (2023); He et al. (2024).
In this paper, we focus on the process of uncovering knowledge that the model has not yet mastered, for supplementation. Throughout each iterative phase of model evolution, we identify the limitations of the model’s knowledge and address them in subsequent updates. This idea is inspired by the ’summarizing mistakes’ learning skill observed in human students. An evaluation is conducted initially to identify a student’s weaknesses in a subject. References related to these errors are then provided to assist the student in mastering the associated knowledge. For instance, if a student is struggling with chemistry, they can review questions answered incorrectly in chemistry tests to pinpoint areas needing improvement and extract relevant knowledge, such as the properties of benzene, from the corresponding chapter in the chemistry textbook. This process enhances his performance in chemistry by identifying and learning from weaknesses or areas of unfamiliarity.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method. The aim is to systematically address the model’s knowledge gaps and rectify incorrect knowledge through continuous evolution. Our contributions are summarized as follows:
  • To endow LLMs with the capacity for continuous iterative evolution, we have drawn inspiration from the ’summarize mistakes’ learning skill observed in human students. We introduce CEM, which explores the knowledge boundaries from LLM mistakes, thereby continuously enhancing the model’s capabilities.
  • To assist LLMs in acquiring and comprehending factual knowledge from supplementary corpora, as well as mitigating knowledge forgetting, we propose two strategies for constructing supplementary corpora.
  • We conduct experiments on three open-source LLMs and two complex knowledge Question-Answering (QA) tasks. The results demonstrate the efficacy of the CEM.
2 Related Work
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al. (2022); Stiennon et al. (2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al. (2022); Bai et al. (2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examples. Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters Hu et al. (2021), prompt tuning Li and Liang (2021), and other techniques.
Continual learning with LLM Continual learning (CL) aims to enable a model to learn from a continuous data stream across time while reducing catastrophic forgetting of previously acquired knowledge Biesialska et al. (2020). With CL, a deployed LLM has the potential to adapt to the changing world without costly re-training from scratch Bubeck et al. (2023). Early works Gururangan et al. (2020); Röttger and Pierrehumbert (2021); Lazaridou et al. (2021); Dhingra et al. (2022) empirically analyze continuing LM pre-training on emerging domain or temporal data, showing the potential to update the base LM with new knowledge. Jang et al. (2022b) explicitly categorize world knowledge as timeinvariant, outdated, and new knowledge, which should be retained, acquired, and updated respectively by an LM when learning continually. Jin et al. (2022); Jang et al. (2022a) additionally implement traditional CL methods to alleviate catastrophic forgetting, a phenomenon in which previously learned knowledge or abilities are degraded due to overwritten parameters Kirkpatrick et al. (2017b).
3 Methods
The objective of this study is to enhance the knowledge base of LLM through continual learning, resulting in iterative improvements in model performance.
3.1 Task Formulation
Question-Answering (QA) tasks are commonly used to evaluate models’ world knowledge and reasoning skills. Our focus is on enhancing the models’ world knowledge rather than their reasoning abilities. These tasks usually involve a series of single-choice questions with multiple options. The ability of a LLM to correctly identify the correct answer within these questions indicates its possession of the relevant knowledge. Conversely, failure to do so suggests a deficiency in the model’s understanding of the knowledge pertinent to the question.
Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i}$
where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplementary corpus $C_{S_{i}}$,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\}$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and retrieve supplementary corpus.
Subsequently, based on the supplementary corpus $C_{S}$, construct the supplementary training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D)$
where $Construct(\cdot)$ is utilized to build an supplementary training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplementary training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplementary training, thereby achieving continual optimization of the model’s performance.
3.2 CEM Method
Specifically, the CEM method proposed in this paper involves three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplementary corpora from multiple data sources, and (3) constructing a supplementary training set and fine-tuning the LLM. Fig.2 shows the process.
Figure 2: The pipeline of CEM method.
Testing LLM and Collecting Incorrectly Answered Questions. We utilize the LLM to answer questions from a QA dataset. Subsequently, by utilizing pattern matching to extract answers from the LLM’s responses, we verify the accuracy of these answers and collect the questions that the LLM has answered incorrectly.
Collecting Supplementary Corpora from Multiple Data Sources. We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies. Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplementary corpora from multiple data sources, including online encyclopedias and search engines.
(1) Search Engine: We submit questions to search engines and use their retrieval techniques to crawl relevant content from the vast corpus of Chinese-language internet. Search engines are characterized by fast response time, high relevance matching, and diverse and rich content. We conduct experiments on two search engines, Baidu and Bing, and optimized our corpus crawling strategy.
Specifically, we search for questions that the LLM answers incorrectly. The search engine ranks results based on relevance. We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplementary corpora. To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer. In this experiment, we utilize Bing as the search engine and set $k=4$, indicating that the main textual content from the detailed pages of the top-4 search results with the highest relevance was retrieved. Fig.5 shows process of collecting corpora from search engine.
(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines. In our experiment, we experiment two online encyclopedias, Baidubaike and Wikipedia.
Specifically, since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval. The main text of the retrieved results serve as supplementary corpora.
In our experiment, we use Wikipedia as the online encyclopedia and instruct ChatGPT-3.5 using a few-shot prompt for entity extraction. Appendix Fig.6 shows process of collecting corpora from online encyclopedia.
Constructing Supplementary Training Set and Fine-Tuning LLM. The Supplementary Training Sets are constructed separately based on the Supplementary Corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction method of the training set will be explained in detail in Section 3.3.
After completing one round of supplementary training, the model is retested on the QA dataset to gather a new set of incorrectly answered questions, which serve as the first step for the next round of supplementary training.
It should be noted that supplementary training should not be endless, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement.
3.3 Supplementary Training Data
Corpus Format Definition
The construction of the supplementary training set involves the following four types of corpus.
(1) Format Corpus: Table 1 shows an example of a Format Corpus. Prior to implementing the CEM method for supplementary training, we fine-tune the LLM using Format Corpora. This equips the LLM with the ability to respond to multiple-choice questions in a standard format, facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions. To maintain this ability, we incorporate a certain proportion of Format Corpus during the supplementary training to prevent the LLM’s response skills from declining.
Example 1: A Case of Format Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 1: An example of a Format Corpus, designed to illustrate its corpus format.
(2) Supplement Corpus: Corpus obtained through multiple data sources.
(3) Extract Corpus: Table 2 shows an example of an Extract Corpus. Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades. We posit that merely providing the LLM with a supplementary corpus containing knowledge is inadequate. Consequently, we designed the Extract Corpus to train the LLM to effectively capture and comprehend knowledge from supplementary corpora, enhancing its utilization of these materials.
Example 2: A Case of Extract Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
(extract prompt)Please answer in the context of the following background material:
(supplement)The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history of the world.
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 2: An example of a Extract Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplementary text has been omitted for the sake of clarity.
(4) General Corpus: Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al. (2021). Therefore, we utilize the training data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
unfiltered
as General Corpus to preserve the model’s general capabilities.
Supplementary Training Set Construction Methods. Through the combination of four types of corpora, we have developed three innovative methods for constructing supplementary training sets. These methods are detailed below.
(1) Normal Method: To mitigate the potential catastrophic forgetting resulting from supplementary training, we mix the General Corpus with the Supplement Corpus, ensuring an equal quantity of those two type of corpus across all construction methods,
$D_{N}=Construct(C_{G},C_{S})$
where $Construct(\cdot)$ is utilized for constructing supplementary training set based on corpora, $C_{G}$ represents the General Corpus, $C_{S}$ indicates the Supplement Corpus, and $D_{N}$ denotes the supplementary training set constructed via the Normal Method.
(2) Extract Method: To evaluate whether the Extract Corpus enhances the LLM’s ability to capture and comprehend knowledge from supplementary corpora as expected, and to associate questions with relevant knowledge when answering, we added the Extract Corpus into $D_{N}$ to generate the supplementary training set $D_{E}$.
$D_{E}=Construct(C_{G},C_{S},C_{E})$
We replaced the Extract Corpus with the General Corpus to obtain the control group $D_{E_{C}}$.
$D_{E_{C}}=Construct(C_{G}^{\prime},C_{S})$
(3) Review Method: Under higher temperature parameter, the responses of the LLM tend to be diverse. When conducting repeat QA tests on the LLM, it may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities. This means that after supplementary training, questions initially answered correctly may be answered incorrectly, resulting in a transition from right to wrong.
Therefore, we strengthen the LLM’s memory of the knowledge implied by correctly answered questions by reviewing the responses provided correctly by the LLM in the previous round.
Specifically, in this round of supplementary training, we sample the correct responses made by the LLM in the previous round and process the correctly answered questions into either ’Format Corpus’ or ’Extract Corpus’ formats. We define it as the Review Corpus, and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we construct $R_{1}$, $R_{2}$, and $R_{3}$,
$C_{R_{1}}=Format(\alpha S_{QA})+Extract((1-\alpha)S_{QA})$
$C_{R_{2}}=Format(S_{QA})$
$C_{R_{3}}=Extract(S_{QA})$
where $\alpha\in(0,1)$, $S_{QA}$ denotes the QAs sampled from the correct responses in the previous round of LLM. $Format(\cdot)$ is employed to process the sampled QAs into Format Corpus format, while $Extract$ is used to process the sampled QAs into Extracted Corpus format.
Note that although Review Corpus looks like a mixture of Extract Corpus and Format Corpus in its compositional form, the corpus of the former comes from the correct responses made by the model and is a proper subset of the test set. In contrast, the corpus used by Extract Corpus and Format Corpus is completely independent of the test set.
We added the Review Corpus into the foundation of $D_{E}$ to generate the supplementary training set $D_{R_{i}}$,
$D_{R_{i}}=Construct(C_{G},C_{S},C_{E},C_{R_{i}}),\quad i=1,2,3$
And we replaced the Review Corpus with the General Corpus to obtain the control group $D_{R_{C}}$.
$D_{R_{C}}=Construct(C_{G}^{\prime\prime},C_{S},C_{E}),$
4 Experiments
4.1 Experimental Settings
Datasets. We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
Xiezhi, a comprehensive evaluation suite designed to assess holistic domain knowledge. It comprises multiple-choice questions across 516 diverse disciplines, ranging from 13 different subjects. Gu et al. (2023).
CMMLU contains about 12,000 questions spanning elementary to advanced professional levels across 67 subjectsLi et al. (2024).
In our experiments, we select 27,265 multiple-choice questions from Xiezhi and 11,917 multiple-choice questions from CMMLU. Appendix Table 7 shows the details of the data division.
Base language models. The objective of our work is to evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even those with smaller parameter scales. We select open-source models with parameter ranges from 6 to 13 billion for our experiment. We use CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, Qwen-7B [^6] https://huggingface.co/Qwen/Qwen-7B-Chat as our instruction fine-tuning base models for CEM. These models have been fine-tuned on the instruction data enabling them to effectively follow instructions. CuteGPT is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023). ChatGLM-6B also is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022). Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
In the instruction fine-tuning stage, CuteGPT-13B and Qwen-7B-chat employ full fine-tuning, and ChatGLM-6B employs Prompt-tuningLiu et al. (2022).
Supplementary training dataset. We construct a variety of supplementary training sets and their respective control groups for the experiments following the method described in Section 3.3. To distinguish between the model using different supplementary training sets, we establish the following naming rules:
(1) Baseline: it represents the model that has not undergone supplementary training and has only been pre-finetuned.
(2) N: N denotes Normal, indicating that the model is supplementary trained using the supplementary training set constructed by the Normal Method.
(3) E and EC: E denotes Extract, indicating that the model is supplementary trained using the supplementary training set constructed by the Extract Method. To obtain its respective control group, Extract Control (EC), we replace the Extract Corpus in the training set with the General Corpus.
(4) R1, R2, R3 and RC: R denotes Review, indicating that the model is supplementary trained using the supplementary training set constructed by the Review Method and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we further name them as R1, R2, and R3. To obtain their respective control group, Review Control (RC), we replace the Review Corpus in the training set with the General Corpus.
In addition, we distinguish between the data sources from which the supplementary corpus is collected and establish the following naming rules:
(1) Wiki: it represents the Supplementary Corpus of the supplementary training set used by the model is from Wikipedia.
(2) Bing: it represents the Supplementary Corpus of the supplementary training set is from Bing.
(3) Mix: it represents the Supplementary Corpus of the supplementary training set is collected from both Bing and Wikipedia and concatenated.
Metrics. The effects of supplementary training will be evaluated using the following metrics:
(1) ACC: it represents the accuracy of the answer of LLM.
(2) W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones in the test set.
(3) R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones in the test set.
In Section 4.2, we combine the above two naming rules to describe the experimental results clearly. For instance, Wiki_E denotes supplementary training using the dataset constructed through the Extract Method. The Supplementary Corpus of this dataset is collected from Wikipedia.
4.2 Results and Analysis
Figure 3: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplementary training set is constructed using the Normal Method.
CEM method enhances the performance of various LLMs across diverse tasks. We conduct tests on LLMs that undergo supplementary training using the CEM method. The training sets are constructed via the Normal Method. The accuracy results are presented in Fig. 3. More complete experimental results are in Appendix Table 4 and 5. After implementing CEM, all LLMs showed significant improvement in accuracy for both QA tasks, demonstrating the effectiveness of CEM in addressing knowledge deficiencies and enhancing LLMs’ performance. For instance, after one round of of CEM optimization with Bing as its supplementary corpus source, the accuracy of Qwen-7B-Chat on the CMMLU task surged by 16.08%.
Smaller LLMs can better benefit from CEM. The effectiveness of LLM tuning with CEM depends on the innate abilities of the LLM and also varies for LLMs with different parameter scales. According to the experimental results presented in Fig. 3, among the open-source models mentioned, ChatGLM-6B, which has the fewest parameters, shows the most significant improvement, up to 13.17% on the Xiezhi task. In contrast, CuteGPT-13B, which has the highest initial answer accuracy, exhibits a smaller improvement of up to 3.63% compared to the smaller LLM.
Extract Corpora and Review Corpora can significantly boost accuracy. We conducted parallel experiments on the supplementary training set construction method described in section 3.3, using CuteGPT-13B on Xiezhi dataset. The results are presented in Fig. 4. CEM with Extract Corpus outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., Mix_E shows a 3.11% increase in accuracy to its control group Mix_EC, and a 10.96% increase compared to the baseline.
Figure 4: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods. ’Avg of Ri’ and ’Best of Ri’ represent the average and best results, respectively, of the three Review Methods.
Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{i=1,2,3}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric. For example, Wiki_R1 shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, Wiki_RC, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the baseline. The results indicate that the W2R is improved by the Extract Corpus. This suggests that the model’s capacity to identify and comprehend the corresponding knowledge in the corpus is enhanced, leading to better error correction. Review Corpus significantly reduced R2W, indicating that the model’s forgetting of correct knowledge was suppressed.
In addition, the three construction methods, $R_{1}$,$R_{2}$ and $R_{3}$, are similar in their ability to enhance LLMs’ performance, which means either form of Review Corpus can be effective. Therefore, if not specified, $R$ will refer to the $R_{1}$ Method in the following.
Supplementary corpora from multiple sources do not always improve the CEM approach more than a single source. The effectiveness of CEM depends on the recall rate when searching corpora and the quality of supplementary corpora, as this determines whether the corpus contains knowledge that the LLM lacks. Moreover, noise and errors within low-quality corpora can negatively impact the LLM. Results in Fig. 4 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated. When using the supplementary training set constructed by Review Method, Mix_R has lower accuracy than Wiki_R and Bing_R despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplementary corpus from different sources. Additionally, an excess of supplementary corpus may dilute the proportion of generalization, excerpt, and comment corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
Metrics Wiki Bing
N E R N E R
Acc Baseline 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R comp. to Baseline 24.46 26.03 28.29 22.07 23.32 27.01
comp. to Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W comp. to Baseline 15.70 13.76 11.29 18.63 17.51 11.61
comp. to Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 3: We employ the identical training data construction method as in the first round to conduct a second round of supplementary training on the LLM. The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
CEM can continuously evolve the LLM through multiple iterations. We conduct a second round of supplementary training on CuteGPT-13B on Xiezhi task using the same dataset construction method and data source as the first round. The results are presented in Table 3. After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round. Both the Extract Method and Review Method are effective in improving the accuracy of the model. For instance, Wiki_R shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to baseline.
Furthermore, the Review Method consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction methods, e.g., the R2W metric relative to Baseline of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E. We argue that the Review Method performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one. This enables the model to be evolved by two or more rounds of supplemental training.
5 Conclusion
In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge. We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplementary corpora from multiple data sources, (3) constructing a supplementary training set and fine-tuning the LLM. Furthermore, we introduce multiple methods for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM. We conduct experiments on several opensource LLMs to validate the efficacy of CEM. The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers. For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
Limitations: This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
In future research, we will further explore three or more rounds of supplementary training to improve the LLM and to mitigate the catastrophic forgetting. In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
Appendix A Process of Collecting Corpora from Multiple Sources
Fig.5 and Fig.6 show the process of obtaining supplementary data from Wikipedia and Bing respectively.
Figure 5: The process of collecting corpora from Bing.
Figure 6: The process of collecting corpora from Wikipedia.
Appendix B Experimental Results of Various LLMs on QA tasks
Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Normal method to construct the supplementary training dataset.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
Table 4: The table shows the experimental results of LLMs using the CEM method on the Xiezhi task.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
Table 5: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
Appendix C Experimental Results with Various Supplementary Training Dataset Construction Methods
Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods.
Data Source Metrics Baseline N E EC R1 R2 R3 RC
Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
W2R - 22.20 24.12 23.42 24.93 24.61 24.82 22.90
R2W - 18.57 15.02 15.65 12.68 12.87 14.35 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.86 54.53 53.90 46.58
W2R - 21.91 23.15 21.37 24.25 25.18 24.99 22.28
R2W - 18.93 17.61 19.18 13.32 13.58 14.02 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.42 54.43 53.37 53.07
W2R - 21.99 25.12 23.67 24.04 25.06 25.00 24.81
R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplementary training set construction methods.
Appendix D Data Division
Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplementary training. The Testing part is used to construct the test set, and the remaining part is used to construct the Format Corpus and Extract Corpus for supplementary training.
Dataset Total Pre-Finetuning Testing Supplementary Training
Xiezhi 27,265 10,000 10,000 7,265
CMMLU 11,917 4,000 5,000 2,917
Table 7: The data division of Xiezhi and CMMLU.