### v3

# Large Language Model Can Continue Evolving From Mistakes

Abstract
As world knowledge evolves and new task paradigms emerge, Large Language Models (LLMs) often fall short of meeting new demands due to knowledge deficiencies and outdated information. Continual Learning (CL) is crucial for keeping LLMs up-to-date and addressing these deficiencies. However, traditional CL approaches struggle to balance task-width generality with task-depth specificity and often lack efficient data collection strategies, leading to increased training costs without addressing the model’s most critical needs.
Inspired by the ‘summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method. This iterative approach continually evaluates LLMs to identify knowledge deficiencies based on their mistakes, collecting relevant data from multiple sources to supplement training in a targeted manner. To enhance the model’s utilization of supplemental knowledge and prevent forgetting, we developed three dataset construction strategies that integrate various types of continual pretraining (CPT) data and continual instruction tuning (CIT) data. Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in LLM accuracy in the best scenarios. Additionally, further experiments confirm the potential of combining CEM with other catastrophic forgetting mitigation strategies, enabling multi-round iterative optimization.
Large Language Model Can Continue Evolving From Mistakes
[…]
1 Introduction
With the exponential growth of data and parameter sizes, coupled with the employment of instruction fine-tuning, Large Language Models (LLMs) have showcased superior performance across a wide array of downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the ongoing emergence of new knowledge and task requirements necessitates continual task-specific adaptation and knowledge supplementation for LLMs. Otherwise, LLMs are likely to generate incorrect outputs based on outdated or lacking knowledge (i.e., ‘hallucinations’) Huang et al. (2023). Accordingly, Continual Learning(CL) is proposed to ensure LLMs’ alignment with the ever-changing world knowledge.
Given the significant training costs, it is more popular to continually pre-train or fine-tune LLMs on a sequence of emerging domain-specific or task-specific data, rather than periodically training them from scratch on the latest corpora (e.g., the most recently archived Wikipedia) Gururangan et al. (2020b). However, common CL tends to be task-width oriented, whereby the model is trained using new task-relevant data to enhance its applicability across a wider range of tasksGururangan et al. (2020b); Cossu et al. (2022), as opposed to focusing on the continual improvement of the model’s performance within a specific domain or task, i.e., task-depth oriented. In addition, building CL datasets is often challenging. Blind data collection fails to ensure access to the high-quality data that the model most requires for task-solving, and may introduce significant training costs due to irrelevant data.
To address these issues, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient method for continually enhancing the performance of LLMs on specific tasks through muiltiple evaluation and supplementation with mistake-relevant knowledge. Our method draws inspiration from the ‘summarizing mistakes’ learning skill of human students. For instance, when a student is struggling with math, they can leverage test mistakes to identify lacking knowledge and extract relevant information from the textbook. Through repeated math tests and summaries, improvements in their math scores have been noted. While recent studies have uncovered successful utilization of mistakes in constructing CoT dataAn et al. (2024), thereby enhancing the model’s performance in mathematical reasoning, there exists potential to extend the leveraging of mistakes to diverse tasks and integrate it within CL paradigms.
Recent studies have shown that the poor performance of LLM on specific tasks may result from two factors: (1) issues beyond its knowledge boundaries, indicating a lack of domain expertise Ren et al. (2023); Zhang et al. (2023b), and (2) knowledge errors within LLM, including those stemming from unreliable pre-training data sources and outdated information Rae et al. (2021); Zhao et al. (2023). We argue that mistakes effectively expose the LLM’s knowledge deficiencies and errors. Addressing these directly represents the most efficient approach to enhancing the model. Therefore, at the start of each CEM round, we test the model and gather cases of incorrect or poor performance. We deconstruct these ‘bad cases’ to identify the most lacking or erroneous knowledge points, then obtain supplemental data from search engines and online encyclopedias to address these gaps. To enhance the LLM’s knowledge acquisition from this supplemental data and prevent forgetting, we devised multiple dataset construction strategies utilizing both unsupervised Continual Pre-training (CPT) data and supervised Continual Instruction Tuning (CIT) dataWu et al. (2024). By integrating CEM with strategies to mitigate catastrophic forgetting, we enable iterative, continual model evolution.
Our contributions are summarized as follows: (1) We propose the CEM method. This method explores knowledge boundaries from LLM’s mistakes, endowing LLMs with the capacity for continual iterative evolution. (2) To assist LLMs in acquiring and comprehending factual knowledge from supplemental corpora, and to mitigate knowledge forgetting, we propose three strategies for constructing supplemental training sets. (3) We conduct experiments on multiple open-source LLMs for two complex knowledge Question-Answering (QA) tasks. Across all scenarios, the CEM method consistently exhibit significant improvements, achieving an optimal accuracy enhancement of 17%. (4) We verify that CEM can be used together with methods for mitigating catastrophic forgetting Li et al. (2022), such as replay-based methods Chaudhry et al. (2019); Wu et al. (2021). This ensures that catastrophic forgetting can be suppressed even with multiple iterations.
2 Related Work
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al. (2022); Stiennon et al. (2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al. (2022); Bai et al. (2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examplesHan et al. (2024). Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters Hu et al. (2022), prompt tuning Li and Liang (2021), and other techniques.
Continual Pre-training (CPT)
Continual pretraining involves regularly updating LLMs with the latest facts or adapting them to specialized domains. To continually pre-train LLMs to update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al. (2019), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022).
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduces a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continuous pretraining for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
Continual Instruction Tuning (CIT)
Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al. (2023a). It can be broadly classified into two categories: 1) task-incremental CIT and 2) domain-incremental CIT.
Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks. A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al. (2023). However, studies have shown this process would cause a catastrophic forgetting of the learned knowledge and problem-solving skills in previous tasks Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tuning of LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
3 Methods
The objective of this work is to enhance the knowledge of LLM through continual learning, resulting in iterative improvements in model performance.
3.1 Task Formulation
Question-Answering (QA) tasks are commonly used to evaluate models’ world knowledge and reasoning skills. Our focus is on enhancing the models’ world knowledge rather than their reasoning abilities. These tasks usually involve a series of single-choice questions with multiple options. The ability of a LLM to correctly identify the correct answer within these questions indicates its possession of the relevant knowledge. Conversely, failure to do so suggests a deficiency in the model’s understanding of the knowledge pertinent to the question.
Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\},$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
3.2 CEM Method
We propose the CEM method, involving three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplemental corpora from multiple data sources, and (3) constructing a supplemental training set and fine-tuning the LLM. Figure 1 shows the process.
Figure 1: The pipeline of CEM method.
Testing LLM and Collecting Incorrectly Answered Questions. We utilize the LLM to answer questions from a QA dataset. Subsequently, by utilizing pattern matching to extract answers from the LLM’s responses, we verify the accuracy of these answers and collect the questions that the LLM has answered incorrectly.
Collecting Supplemental Corpora from Multiple Data Sources. We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies. Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplemental corpora from multiple data sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are characterized by fast response time, high relevance matching, and diverse and rich content. We search for questions that the LLM answers incorrectly. The search engine ranks results based on relevance. We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplemental corpora. To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer. Additionally, we optimize the strategy for crawling the corpus to filter out irrelevant information and spam, such as advertisements.
In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 4 in Appendix shows a case of corpus collection from Bing.
(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines. Since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval. The main text of the retrieved results serve as supplemental corpora.
In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction. Figure 5 in Appendix shows a case of corpus collection from Wikipedia.
Constructing Supplemental Training Set and Fine-Tuning LLM. The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction method of the training set will be explained in detail in Section 3.3.
After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions. This can serve as the first step for the next round of supplemental training.
It is important to note that supplemental training should not be never-ending, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement. In the following experiments, we will examine how the number of supplemental training rounds affects model improvement.
3.3 Supplemental Training Data
Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
(1) Supplemental Corpus($C_{\underline{S}}$): Supplemental Corpus obtained from multiple data sources contains background knowledge that the LLM lacks and is used similarly to continuous pre-training Gupta et al. (2023).
Example 1: A Case of Normative Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 1: A case of a Normative Corpus, designed to illustrate its corpus format.
Example 2: A Case of Extractive Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
(extract prompt)Please answer in the context of the following background material:
(supplement)The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history of the world.
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 2: A case of a Extractive Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplemental text has been omitted for clarity.
(2) Normative Corpus($C_{\underline{N}}$): To equip the LLM with the ability to respond to multiple-choice questions in a normative format, we sample multiple-choice questions that are independent of the test set and process them into the format shown in Table 1 to build Normative Corpus. Prior to supplemental training, we finetune the LLM using Normative Corpus(i.e. pre-finetuning), facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions. To maintain this ability, we incorporate a small proportion of Normative Corpus during the supplemental training to prevent the LLM’s response skills from declining.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To improve the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample multiple-choice questions that are independent of the test set and the Normative Corpus. These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
(4) Review Corpus($C_{\underline{R}}$): The responses of the LLM tend to be diverse under high temperature Caccia et al. (2018). During repeated QA testing, LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting the accuracy.
Therefore, to strengthen the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly of the LLM in the previous round and process some of them in Normative Corpus format and the rest in Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus can be described as:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round of LLM. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al. (2021). Therefore, to preserve the model’s general capabilities on other tasks, we use the data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
unfiltered
as General Corpus.
Supplemental Training Set Construction Strategy. We have developed three innovative strategies for constructing supplemental training sets by combining five types of corpus. The strategies are detailed below.
(1) Plain Strategy: To mitigate the potential catastrophic forgetting and to maintain the LLM’s ability to provide normative responses, we mix three types of corpus to construct the supplemental training set:
$D_{P}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges various types of corpus together and $D_{P}$ denotes the supplemental training set constructed via the Plain Strategy.
(2) Extractive Strategy: To enhance the LLM’s ability to capture and comprehend knowledge from supplemental material, we integrate the Extractive Corpus:
$D_{E}=Concat(C_{S},C_{N},C_{G},C_{E}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ denotes the supplemental training set constructed via the Extractive Strategy.
(3) Review Strategy: To enhance the model’s retention of acquired knowledge and prevent correct responses from being altered to incorrect ones over multiple iterations, we integrate the Review Corpus:
$D_{R}=Concat(C_{S},C_{N},C_{G},C_{E},C_{R}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ denotes the supplemental training set constructed via the Review Strategy.
4 Experimental Setup
4.1 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
Xiezhi, a comprehensive evaluation suite designed to assess holistic domain knowledge. It comprises multiple-choice questions across 516 diverse disciplines, ranging from 13 different subjects. Gu et al. (2023).
CMMLU contains about 12,000 questions spanning elementary to advanced professional levels across 67 subjectsLi et al. (2024).
In our experiments, we select 27,265 multiple-choice questions from Xiezhi and 11,917 multiple-choice questions from CMMLU. Appendix Table 7 shows the details of the data division.
4.2 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion. Our instruction fine-tuning base models include CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, and Qwen-7B-Chat [^6] https://huggingface.co/Qwen/Qwen-7B-Chat. These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
CuteGPT-13B is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023).
ChatGLM-6B is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022).
Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
4.3 Metrics
The effectiveness of supplemental training will be assessed using the following metrics:
ACC: it represents the accuracy of the answer of LLM.
W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
5 Experiments
We construct various supplemental training sets and their corresponding control groups for the experiments, following the method described in Section 3.3. To differentiate between models trained on different supplemental training sets, we establish the following naming conventions:
INIT: it represents the model, which has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
CEM-N: N denotes Normative, representing the model that has undergone supplemental training using the training set constructed by the Plain Strategy.
[…]
We append suffixes to the aforementioned naming conventions to differentiate the data sources of the supplemental corpus. These suffixes include: ‘Wiki’, ‘Bing’ and ‘Mix’ where ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated. For instance, ‘E_Wiki’ denotes the model supplemented trained using the dataset constructed through the Extractive Strategy. The Supplemental Corpus of this dataset is collected from Wikipedia.
5.1 Main Results
We conduct experiments to analyze the accuracy after applying CEM on the test data. The results are shown in Figure 2. Our observations are summarized as follows:
Figure 2: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplemental training set is constructed using the Plain Strategy.
Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two given tasks. This implies that despite the high availability of supplemental data and the small size of the training set, the CEM method effectively complements the model’s knowledge deficiencies. For instance, the ChatGLM-6B model shows an improvement of up to 16.08% on the CMMLU task.
Secondly, the effectiveness of tuning LLMs with CEM depends on their innate abilities and also varies across LLMs with different parameter scales. Among the mentioned open-source models, ChatGLM-6B, which has the fewest parameters, exhibits the most significant improvement, up to 13.17% on the Xiezhi task. In contrast, CuteGPT-13B, with the highest initial answer accuracy and the most parameters, shows a smaller improvement of up to 3.63% compared to the smaller LLM.
Construction Strategy Analysis. To investigate the effectiveness of each component in supplemental training set, we conduct experiments on LLMs trained using various dataset construction strategies, the results are shown in Figure 3. Our observations are summarized as follows:
Figure 3: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies. ’Avg of $R_{\alpha}$’ and ’Best of $R_{\alpha}$’ represent the average and best results, respectively, of the Review Strategy with different $\alpha$.
Firstly, Extractive Corpus and Review Corpus can significantly boost LLMs’ performance. CEM with Extractive Corpus ($C_{E}$) outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., E_Mix shows a 3.11% increase in accuracy to its control group E’_Mix, and a 10.96% increase compared to INIT. Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{\alpha=0,0.5,1}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric. For example, R0.5_Wiki shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, R’_Wiki, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the INIT.
The results indicate an improvement in W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the corpus. This enhancement leads to more effective error correction. Additionally, the Review Corpus significantly reduced R2W, indicating suppression of the model’s tendency to forget correct knowledge.
Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source. The effectiveness of CEM depends on the recall rate when searching corpus and the quality of supplemental corpus, as this determines whether the corpus contains knowledge that the LLM lacks. Moreover, noise and errors within low-quality corpora can negatively impact the LLM. Results in Figure 3 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated. When using the supplemental training set constructed by Review Strategy, R_Mix has lower accuracy than R_Wiki and R_Bing despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplemental corpus from different sources. Additionally, an excess of supplemental corpus may dilute the proportion of General, Extractive, and Normative Corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of the Review Corpus. According to Table 6 in Appendix, R0, R0.5 and R1 are similar in their ability to enhance LLMs’ performance. This may be because the models themselves have already mastered the knowledge, and replaying it in either format serves to consolidate that knowledge. Therefore, if not specified, R will refer to the R0.5 Strategy in the following.
5.2 Multiple Iterations Analysis
Metrics Wiki Bing
CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 3: We employ the identical training data construction strategy as in the first round to conduct a second round of supplementary training on the LLM. The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
We conduct a second round of supplemental training on CuteGPT-13B on Xiezhi task using the same dataset construction strategy and data source as the first round. The results are presented in Table 3. After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round. Both the Extractive Strategy and Review Strategy are effective in improving the accuracy of the model. For instance, R_Wiki shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to INIT.
Furthermore, the Review Strategy consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction strategies, e.g., the R2W metric relative to INIT of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E. We argue that the Review Strategy performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one. This enables the model to be evolved by two or more rounds of supplemental training.
6 Discussion
Why can the Extractive Corpus be effective? We observe that models employing the Extractive Corpus exhibit higher accuracy and W2R metric. Unlike the Review Corpus, where improving model performance is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, whereas the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions may be a structured induction of supplemental material, where the Extractive Corpus explicitly links complex information within the corpus to the resulting induction (i.e., the question and options). Questions and options can act as cues and guides for the model to gather relevant information from the corpus, considering how this knowledge is accessed through questions. This, in turn, can enhance the model’s ability to extract useful information from the corpus while eliminating redundant and irrelevant information.
7 Conclusion
In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge. We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplemental corpora from multiple data sources, (3) constructing a supplementl training set and fine-tuning the LLM. Furthermore, we introduce multiple strategies for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers. For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
8 Limitations
This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
In future research, we will further explore three or more rounds of supplemental training to improve the LLM and to mitigate the catastrophic forgetting. In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
Appendix A Process of Collecting Corpora from Multiple Sources
Fig.4 and Fig.5 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
Figure 4: The process of collecting corpora from Bing.
Figure 5: The process of collecting corpora from Wikipedia.
Appendix B Experimental Results of Various LLMs on QA tasks
Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
Table 4: The table shows the experimental results of LLMs using the CEM method on the Xiezhi task.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
Table 5: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
W2R - 22.20 24.12 23.42 24.93 24.61 24.82 22.90
R2W - 18.57 15.02 15.65 12.68 12.87 14.35 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.86 54.53 53.90 46.58
W2R - 21.91 23.15 21.37 24.25 25.18 24.99 22.28
R2W - 18.93 17.61 19.18 13.32 13.58 14.02 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.42 54.43 53.37 53.07
W2R - 21.99 25.12 23.67 24.04 25.06 25.00 24.81
R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
Appendix D Data Division
Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing supplemental Training
Xiezhi 27,265 10,000 10,000 7,265
CMMLU 11,917 4,000 5,000 2,917
Table 7: The data division of Xiezhi and CMMLU.