### v4

# Large Language Model Can Continue Evolving From Mistakes

Abstract
As world knowledge evolves and new task paradigms emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings. In practical applications, LLMs often require both continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new task paradigms and acquire necessary knowledge for task-solving. However, it remains challenging to collect CPT data that addresses the knowledge deficiencies in models while maintaining adequate volume, and improving the efficiency of utilizing this data also presents significant difficulties. Inspired by the ’summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient approach for collecting CPT data and continually improving LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To efficiently utilize these CPT data and mitigate forgetting, we design a novel CL training set construction paradigm that integrates parallel CIT and CPT data. Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in accuracy in the best case. Furthermore, additional experiments confirm the potential of combining CEM with catastrophic forgetting mitigation methods, enabling iterative and continual model evolution.
Large Language Model Can Continue Evolving From Mistakes
Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
1 Introduction
With the exponential growth of data and parameter sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and task requirements necessitates ongoing task adaptation and knowledge supplementation for LLMs. Otherwise, LLMs may fall into hallucinations and generate incorrect responses Huang et al. (2023). Therefore, Continual Learning (CL) is proposed to ensure LLMs’ alignment with the evolving world knowledge and demands.
Figure 1: The figure illustrates two potential triggers for poor model performance: (1) Task Paradigm Unfamiliarity, and (2) Lack of task-relevant knowledge. Unfamiliarity with the task paradigm can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations. Instruction tuning addresses only the former.
Wu et al. (2024) categorize Continual Learning into Continual Pre-training (CPT) and Continual Instruction Tuning (CIT). CPT requires a larger volume of training data to enrich LLMs’ knowledge for new domains, while CIT continually fine-tunes LLMs on emerging task-specific instruction data to adapt the models to evolving task paradigms with much smaller training costs and data requirements Razdaibiedina et al. (2023); Gururangan et al. (2020b); Cossu et al. (2022). However, recent work has highlighted the limitations of instruction tuning, suggesting it is merely a superficial alignment process focused on learning interaction styles and LLMs primarily acquire knowledge during the pre-training stage Zhou et al. (2023). This will trigger the bad response shown in Figure 1, where LLMs’ poor performance may result from a lack of requisite knowledge to solve the task rather than unfamiliarity with the task paradigm Ren et al. (2023); Zhang et al. (2023b). CPT is more effective than CIT in addressing this issue.
However, using CPT to supplement the model with the lacking knowledge required for task-solving presents several issues: First, collecting CPT data is challenging, as it is difficult to a priori identify the lacking knowledge while ensuring a sufficiently large data scale. Blind data collection and insufficient data volume may introduce irrelevant information, weaken generalization, and pose overfitting risks Zhang et al. (2024); Gekhman et al. (2024). Second, due to the weak constraints of self-supervised CPT data during parameters updates and the scale gap compared to initial pre-training data, efficiently utilizing CPT data and avoiding catastrophic forgetting remain difficult.
To address these issues, recent work has explored converting CPT data into CIT data, such as transforming it into reading comprehension questions to avoid the problem of insufficient data volume. Cheng et al. (2024) or using highly relevant QA pairs for CIT before CPT to enhance the data utilization efficiency Jiang et al. (2024). However, these methods demand high-quality data and the assistance of oracle LLM, and they overlook the crucial challenge of CPT mentioned above. Inspired by the ‘summarizing mistakes’ learning skill of human students, we argue that a model’s mistakes in ’exams’ reflect its inherent knowledge deficiencies. CPT data that directly addresses these gaps is the most effective for supplementing the model’s knowledge. Currently, the successful utilization of mistakes in constructing CoT data An et al. (2024); Tong et al. (2024) has enhanced the model’s performance in mathematics and reasoning, indicating the potential to leverage mistakes in continual learning.
In this paper, we propose the Continue Evolving from Mistakes(CEM) method, aiming to provide a data-efficient method for collecting CPT data and continually enhancing the performance of LLMs through multiple evaluation and supplementation with mistake-relevant knowledge. Specifically, at the start of each CEM round, we test the model, identify cases of poor performance, and deconstruct them to pinpoint knowledge deficiencies. We then gather abundant supplemental data based on mistake analysis from search engines and online encyclopedias as CPT data to address these gaps. To efficiently utilize these CPT data and mitigate forgetting, we designed a novel CL training set construction paradigm that integrates parallel CIT and CPT data. By integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution.
Our contributions are summarized as follows: (1) We propose the CEM method, which identifies knowledge deficiencies from LLMs’ mistakes and collects abundant relevant CPT data, enabling continual iterative evolution. (2) To assist LLMs in acquiring and comprehending factual knowledge from CPT data and mitigating knowledge forgetting, we propose three strategies for constructing supplemental training set using a parallel paradigm with CIT and CPT data. (3) We conduct experiments on multiple open-source LLMs for two complex knowledge QA tasks. The CEM method consistently demonstrates significant improvements across all scenarios, , achieving up to a 17.00% accuracy enhancement. (4) We verify that CEM can be combined with methods to mitigate catastrophic forgetting, such as replay-based methods Chaudhry et al. (2019); Wu et al. (2021), ensuring suppression of catastrophic forgetting over multiple iterations.
2 Related Work
Continual Pre-training (CPT)
Continual pre-training involves regularly updating LLMs with the latest facts or adapting them to specialized domains Wu et al. (2024). To continually pre-train LLMs and update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al. (2019), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022).
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduced a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continual pre-training for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
Continual Instruction Tuning (CIT)
Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al. (2023a). It can be broadly classified into two categories: (1) task-incremental CIT and (2) domain-incremental CIT Wu et al. (2024).
Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks. A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al. (2023). However, studies have shown this process causes catastrophic forgetting of the learned knowledge and problem-solving skills from previous tasks Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
3 Methods
3.1 Task Formulation
Question Answering (QA) tasks are commonly used to evaluate LLM’s domain knowledge. These tasks typically involve a series of single-choice questions with multiple options. If an LLM correctly answers these questions, it indicates a strong grasp of the relevant knowledge. Conversely, incorrect answers reveal deficiencies in understanding.
Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$ as CPT data,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\ldots\},$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
3.2 CEM Method
We propose the CEM method, as illustrated in Figure 2. The main steps are as follows:
Figure 2: The pipeline of CEM method.
Step 1&2: Testing LLM and Collecting Incorrectly Answered Questions. We use the LLM to answer questions from a QA dataset. By employing pattern matching to extract answers from the LLM’s responses, we calculate the accuracy and collect the questions, along with their options, that the LLM answered incorrectly.
Step 3: Obtaining Supplemental Corpus from Multiple Data Sources. Based on the implicit information of knowledge deficiencies contained within questions incorrectly answered by the LLM, we collect supplemental corpus from multiple data sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are characterized by high relevance matching and diverse content. We search for questions the LLM answers incorrectly. The search engine ranks results by relevance, and we select the top-k results to access their detailed pages, extracting the main textual content as supplemental corpus. To prevent the LLM from merely memorizing answers, we filter out results that contain both the original question and its answer. Additionally, we optimize the corpus crawling strategy to exclude irrelevant information and spam, such as advertisements.
In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 6 in Appendix shows the process of corpus collection from Bing.
(2) Online Encyclopedia: Compared to search engines, online encyclopedias provide more credible and professional results, despite lower hit rates. Since encyclopedias are typically organized by keyword entries, we conduct entity extraction from the questions and options, using the extracted entity keywords for retrieval. The main text of the retrieved results serves as the supplemental corpus.
In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction. Figure 7 in Appendix shows the process of corpus collection from Wikipedia.
Step 4&5: Constructing Supplemental Training Set and Fine-Tuning LLM. The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction strategies of the training set will be explained in detail in Section 3.3.
After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions. This can serve as the first step for the next round of CEM.
It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement. Subsequent experiments explore the impact of the number of training rounds on model performance.
3.3 Supplemental Training Data
Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
(1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data.
Below are the three types of CIT data:
(2) Normative Corpus($C_{\underline{N}}$): To enable the LLM to answer multiple-choice questions in a normative format, we sample questions independent of the test set and process them into the format shown in Table 1 to build the Normative Corpus. Prior to supplemental training, we finetune the LLM using the Normative Corpus (i.e., pre-finetuning), which facilitates extracting answers from the LLM’s responses to calculate initial performance and allows for the collection of incorrectly answered questions. To maintain this ability, we incorporate a small proportion of the Normative Corpus during supplemental training.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions independent of the test set and the Normative Corpus. These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
(4) Review Corpus($C_{\underline{R}}$):
The responses of the LLM tend to be diverse under high temperature settings Caccia et al. (2018). During repeated QA testing, the LLM may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round. We process some of these responses in the Normative Corpus format and the rest in the Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus is described as follows:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered as the General Corpus to preserve the model’s general capabilities on other tasks.
Supplemental Training Set Construction Strategy. We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data. Specifically, we construct three types of supplemental training sets by combining five types of corpora. The following sections detail these construction strategies.
(1) Plain Strategy:
$D_{\underline{P}}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges these corpora, and $D_{P}$ represents the supplemental training set constructed using the Plain Strategy.
(2) Extractive Strategy:
$D_{\underline{E}}=Concat(C_{S},C_{N},C_{G},{\color[rgb]{1,0,0}C_{E}}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ represents the supplemental training set constructed using the Extractive Strategy.
(3) Review Strategy:
$D_{\underline{R}}=Concat(C_{S},C_{N},C_{G},C_{E},{\color[rgb]{1,0,0}C_{R}}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ represents the supplemental training set constructed using the Review Strategy.
4 Experimental Setup
4.1 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024).
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities. Appendix Table 6 shows the details of the data division and usage.
4.2 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion. Our instruction fine-tuning base models include CuteGPT-13B [^2] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al. (2023), ChatGLM-6B [^3] https://huggingface.co/THUDM/chatglm3-6b Du et al. (2022), and Qwen-7B-Chat [^4] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al. (2023). These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
4.3 Metrics
The effectiveness of supplemental training will be assessed using the following metrics in the main experiments:
ACC: it represents the accuracy of the answer of LLM.
W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
To analyze the potential performance degradation caused by catastrophic forgetting, we also use the following metrics:
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{i}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of taks $i$ after applying CEM on task $k$.
5 Experiments
We construct supplemental training sets and control groups for the experiments, following the strategies described in Section 3.3. To distinguish between models trained on different supplemental training sets, we use the following naming conventions:
INIT represents the model that has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
CEM-P: P denotes Plain, representing the model trained with the supplemental training set constructed by the Plain Strategy.
[…]
We append suffixes to indicate the data sources of the Supplemental Corpus: ‘Wiki’, ‘Bing’ and ‘Mix’. ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated. For example, ‘E-Wiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
5.1 Main Results
We used the Plain Strategy to construct supplementary training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM. The main results are shown in Figure 3. Our observations are summarized as follows:
Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two tasks. This indicates that the CEM method effectively addresses the model’s knowledge deficiencies, even with the small training set and readily available supplemental data. For instance, the ChatGLM-6B improves by up to 16.08% on the CMMLU task.
Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs. For example, on the Xiezhi task, ChatGLM-6B, with the fewest parameters, improves by up to 13.17%. In contrast, CuteGPT-13B, with the highest initial accuracy and most parameters, shows a smaller improvement of up to 3.63%.
Construction Strategy Analysis. To investigate the effectiveness of each component in the supplemental training set, we conducted experiments on CuteGPT-13B with various dataset construction strategies. We set the volume of both Extractive Corpus and Review Corpus to 1,000. The results are shown in Figure 4. Below is a summary of our observations:
[…]
Firstly, Extractive and Review Corpus significantly enhance LLM performance. CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics. Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric. The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source. Results in Figure 4 show no significant difference in CEM effectiveness whether using Wikipedia, Bing, or a combination of both. Indeed, R-Mix has lower accuracy than R-Wiki and R-Bing despite having more data. This may be due to redundancy from overlapping sources. Additionally, an excess of supplemental corpus can dilute the proportion of CIT data, reducing their effectiveness in mitigating catastrophic forgetting and enhancing knowledge extraction, and increasing the risk of overfitting.
Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of Review Corpus. Table 5 in Appendix shows that R0, R0.5, and R1 have similar effects on LLM performance. This suggests that reviewing the knowledge in any format consolidates it. Therefore, unless specified, R will refer to the R0.5 in the following.
5.2 Multiple Iterations Analysis
We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, using the same training set construction strategy and data source as the first round. We select CIT data independent of the first round but maintain the same data volume. The experimental results presented in Appendix Table 8 show that after the second round of CEM, the accuracy of the LLM’s answers significantly improved across all dataset construction strategies. In the best case, R-Wiki shows a 4.75% improvement in accuracy compared to its previous round and a 17.00% improvement compared to INIT.
Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies. For example, the R2W metric of R-Wiki decreases by 7.02% compared to P-Wiki and by 5.90% compared to E-Wiki. We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplementary training.
Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM combined with the random replay method. Here, ER indicates the model’s enhancement rate on the Xiezhi task, while AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
Catastrophic Forgetting Analysis
To analyze catastrophic forgetting that may arise from iterations, we use CuteGPT-13B and test it on Xiezhi, HotpotQA, and GSM8K tasks to evaluate the model’s domain knowledge, reasoning, and mathematical abilities. We employ the CEM method to enhance performance on Xiezhi and assess any performance declines on the other two tasks. Initially, we sample 2,000 instances from each of the three datasets for pre-finetuning, then test to establish baseline. We use a method of random replay to mitigate catastrophic forgetting. Specifically, in each training round, we randomly replay 500 instances from both HotpotQA and GSM8k to replace the original General Corpus.
Figure 5 shows that even after three rounds of CEM, performance on Xiezhi still exhibits significant improvement and maintains an upward trend. Additionally, the combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70% in most cases. However, we observe that CEM-R, with its larger data capacity, achieves significant performance improvements but also shows the greatest decline on other tasks. This may be attributed to the increased volume of CPT data diluting the effect of the replay data. Combining CEM with other advanced replay-based CL methods warrants further investigation.
6 Discussion
Why can the Extractive Corpus be effective? LLMs using the Extractive Corpus show higher accuracy and W2R metrics. Unlike the Review Corpus, where the role in performance improvement is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, while the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions act as a structured induction of the supplemental material, where the Extractive Corpus explicitly links complex information to the resulting induction (i.e., the question and options). Questions and options can serve as cues and guides for the model to gather relevant information from the corpus, helping it consider how this knowledge is accessed through questions. This process enhances the model’s ability to extract useful information from the corpus while filtering out redundant and irrelevant details.
7 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement. To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations. In the best case, accuracy enhancement reaches up to 17.00%.
Appendix A Format of CIT Data
Table 1 and Table 2 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
An Example of Normative Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 1: An example illustrating the format of the Normative Corpus.
An Example of Extractive Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
(extract prompt)Please answer in the context of the following background material:
(supplement)The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history of the world.
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 2: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix B Process of Collecting Corpora from Multiple Sources
Fig.6 and Fig.7 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
Figure 6: The process of collecting supplementary corpus from Bing.
Figure 7: The process of collecting supplementary corpus from Wikipedia.
Appendix C Experimental Results of LLMs on QA tasks
Table 3 and Table 4 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
Table 3: The table shows the experimental results of LLMs using the CEM method on the Xiezhi task.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
Table 4: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
Table 5 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
W2R - 22.20 24.12 23.42 24.93 24.61 24.82 22.90
R2W - 18.57 15.02 15.65 12.68 12.87 14.35 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.86 54.53 53.90 46.58
W2R - 21.91 23.15 21.37 24.25 25.18 24.99 22.28
R2W - 18.93 17.61 19.18 13.32 13.58 14.02 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.42 54.43 53.37 53.07
W2R - 21.99 25.12 23.67 24.04 25.06 25.00 24.81
R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
Table 5: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
Appendix E Data Division
Table 6 and Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing Supplemental Training
Xiezhi 27,265 10,000 10,000 7,265
CMMLU 11,917 4,000 5,000 2,917
Table 6: The data division and usage of Xiezhi and CMMLU.
Dataset Total Pre-Finetuning Testing Random Replay
HotpotQA 7,000 2,000 5,000 500
GSM8K 7,000 2,000 5,000 500
Table 7: The data division and usage of HotpotQA and GSM8K, where the data used for Random Replay is included in pre-finetuning.
Appendix F Experimental Results of Multiple Rounds of CEM
We employ the identical training data construction strategy as in the first round to conduct a second round of CEM on the LLM. Table 8 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after two rounds of CEM.
Metrics Wiki Bing
CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 8: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task, where "v.s. INIT" indicates W2R and R2W calculated based on the initial evaluation results, and "v.s. Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
Appendix G Experimental Results of Catastrophic Forgetting Analysis
Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
Round Strategy Task Accuracy Relative Rate
Xiezhi GSM8K HotpotQA ER AFR
INIT 40.84 13.20 80.60 1.00 1.00
Round 1 CEM-P 42.02 6.64 51.60 1.03 0.730
CEM-E 42.76 8.36 55.52 1.05 0.681
CEM-R 51.49 7.54 55.62 1.26 0.673
Round 2 CEM-P 48.68 9.94 57.84 1.19 0.723
CEM-E 49.47 9.76 60.24 1.21 0.746
CEM-R 54.23 6.88 56.68 1.33 0.703
Round 3 CEM-P 50.58 10.02 60.96 1.24 0.757
CEM-E 52.69 9.46 62.98 1.29 0.772
CEM-R 58.91 7.48 59.86 1.44 0.678
Table 9: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi, GSM8K, and HotpotQA tasks.