### v5

# Large Language Model Can Continue Evolving From Mistakes

Abstract
As world knowledge evolves and new task schemas emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings. LLMs typically require continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new tasks and acquire essential knowledge. However, collecting sufficient CPT data while addressing knowledge gaps remains challenging, as does optimizing the efficiency of utilizing this data. Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To enhance data utilization and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT data. Experiments demonstrate that CEM significantly enhances model performance and continual evolution. The code and dataset are available in the GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
Large Language Model Can Continue Evolving From Mistakes
Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
1 Introduction
With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs. Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al. (2023). Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial. Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al. (2024).
Figure 1: Two potential triggers for poor model performance: (1) Task Schema Unfamiliarity, and (2) Lack of Task-relevant Knowledge. Unfamiliarity with the task schema can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations. Instruction tuning has been shown to be effective for addressing the former, but poor for the latter Gekhman et al. (2024a); Zhou et al. (2023).
Despite CIT’s smaller data size requirement and lower risk of overfitting compared to CPT, it is limited in new knowledge injection, as Zhou et al. (2023) highlight that instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and does not effectively help the model acquire new knowledge. Thus, it cannot address all the issues shown in Figure 1, where LLMs’ poor performance may stem from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al. (2023); Zhang et al. (2023b). In contrast, CPT is more effective in addressing the lack of knowledge.
However, using CPT to supplement the model with lacking knowledge still presents several challenges. First, collecting CPT data is challenging as identifying the lacking knowledge a priori while ensuring a sufficiently large data size is difficult. Blind data collection and insufficient data volume may introduce irrelevant or redundant information, weaken generalization, and increase overfitting risks Zhang et al. (2024); Gekhman et al. (2024b). Second, efficiently utilizing CPT data and mitigating degradation are difficult due to weak constraints on parameter updates, leading to larger parameter shifts and a data size gap compared to the initial pre-training.
To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al. (2022); Jin et al. (2022). However, these methods lack focus and efficiency and fail to dynamically update model deficiencies. Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al. (2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al. (2024) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data. Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation. To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’. We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies. Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes for directly address these deficiencies. To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data. In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution. The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 17.00% accuracy enhancement.
2 Related Work
Continual Pre-training (CPT)
Continual pre-training involves regularly updating LLMs with the latest facts or adapting them to specialized domains Wu et al. (2024). To continually pre-train LLMs and update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al. (2019), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022).
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduced a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continual pre-training for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
Continual Instruction Tuning (CIT)
Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al. (2023a). It can be broadly classified into two categories: (1) task-incremental CIT and (2) domain-incremental CIT Wu et al. (2024).
Task-incremental Continual Instruction Tuning (CIT) aims to continually fine-tune LLMs on new task-specific instructions. While a straightforward approach involves generating new instruction-tuning data and fine-tuning LLMs directly Wang et al. (2023), this method can lead to catastrophic forgetting of previously learned knowledge and skills Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
3 Methods
3.1 Task Formulation
Question Answering (QA) tasks are commonly used to evaluate LLM’s domain knowledge. These tasks typically involve a series of single-choice questions with multiple options. If an LLM correctly answers these questions, it indicates a strong grasp of the relevant knowledge. Conversely, incorrect answers reveal deficiencies in understanding.
Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$ as CPT data,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\ldots\},$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
3.2 CEM Method
We propose the CEM method, as illustrated in Figure 2. The main steps are as follows:
Figure 2: The pipeline of CEM method.
Step 1&2: Testing LLM and Collecting Incorrectly Answered Questions. We use the LLM to answer questions from a QA dataset. By employing pattern matching to extract answers from the LLM’s responses, we calculate the accuracy and collect the questions, along with their options, that the LLM answered incorrectly.
Step 3: Obtaining Supplemental Corpus from Multiple Data Sources. We first employ the lightweight LLM, Qwen2-7B-instruct [^2] https://huggingface.co/Qwen/Qwen2-7B-Instruct Yang et al. (2024), to identify knowledge points involved in questions erroneously answered by the tested LLM. For further details, see Appendix A. The extracted knowledge points are then utilized as keywords to aggregate supplemental corpora from diverse sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are known for high relevance matching and diverse content. We search for extracted knowledge points using Bing and select the top-k results based on relevance, scraping the main textual content as a supplemental corpus. To prevent LLMs from simply memorizing answers, we filter out results that contain both the original question and its answer. Moreover, we use newspaper3k [^3] https://github.com/codelucas/newspaper and readability [^4] https://github.com/buriy/python-readability to exclude irrelevant and spam content.
(2) Online Encyclopedia: Compared to search engines, online encyclopedias offer more credible and professional results, despite having lower hit rates. Since encyclopedias are typically organized by keyword entries, we employ the state-of-the-art massive text embedding model, Conan-embedding-v1555https://huggingface.co/TencentBAC/Conan-embedding-v1 Li et al. (2024b), to generate vector embeddings for Wikipedia entry titles. FAISS [^6] https://github.com/facebookresearch/faiss Douze et al. (2024) is then used to build an efficient vector index, enabling searches based on inner product similarity. By calculating the similarity between input keyword embeddings and the FAISS-indexed entry embeddings, we retrieve the top-k results that exceed a predefined similarity threshold, serving as the supplemental corpus.
Appendix C presents the experimental parameters and processes.
Step 4&5: Constructing Supplemental Training Set and Fine-Tuning LLM. The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction strategies of the training set will be explained in detail in Section 3.3.
After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions. This can serve as the first step for the next round of CEM.
It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement. Subsequent experiments explore the impact of the number of training rounds on model performance.
3.3 Supplemental Training Data
Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
(1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data. We employ Qwen2-72B-Instruct [^7] https://huggingface.co/Qwen/Qwen2-72B-Instruct Yang et al. (2024) for entailment checking to verify whether the Supplemental Corpus contains information that helps the model answer correctly. Details are provided in the Appendix B.
Below are the three types of CIT data:
(2) Normative Corpus($C_{\underline{N}}$):
To enhance the LLM’s understanding of task schemas and formats, task-relevant instruction fine-tuning is commonly applied. For QA tasks, we sample questions from the training set and format them as shown in Table 3 to build the Normative Corpus, enabling the LLM to answer multiple-choice questions in a standardized manner. Before supplemental training, we fine-tune the LLM using the Normative Corpus (i.e., pre-finetuning). A small portion of the Normative Corpus is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions from the training set and format them as shown in Table 4 to build the Extractive Corpus. We employ Qwen2-72B-Instruct for entailment checking to filter the corpus. A minimal amount of Extractive Corpus data yields effective results, ensuring manageable inference costs. Details are provided in the Appendix E.
(4) Review Corpus($C_{\underline{R}}$):
The responses of the LLM tend to be diverse under high temperature settings Caccia et al. (2018). During repeated QA testing, the LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round. We process some of these responses in the Normative Corpus format and the rest in the Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus is described as follows:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^8] https://huggingface.co/datasets/Aeala/ShareGPT_
Vicuna_unfiltered
as the General Corpus to preserve the model’s general capabilities on other tasks and instruction-following ability.
Supplemental Training Set Construction Strategy. We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data. Specifically, we construct three types of supplemental training sets by combining five types of corpora. The following sections detail these construction strategies.
(1) Plain Strategy:
$D_{\underline{P}}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges these corpora, and $D_{P}$ represents the supplemental training set constructed using the Plain Strategy.
(2) Extractive Strategy:
$D_{\underline{E}}=Concat(C_{S},C_{N},C_{G},{\color[rgb]{1,0,0}C_{E}}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ represents the supplemental training set constructed using the Extractive Strategy.
(3) Review Strategy:
$D_{\underline{R}}=Concat(C_{S},C_{N},C_{G},C_{E},{\color[rgb]{1,0,0}C_{R}}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ represents the supplemental training set constructed using the Review Strategy.
4 Experimental Setup
4.1 Training Setup
We train the models on 2 A800 GPUs using ZeRO (Rajbhandari et al., 2020) stage 1 from DeepSpeed (Rasley et al., 2020). We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 16 with a maximum sequence length of 2,048 and 1 training epoch.
We explore the impact of full fine-tuning, LoRA fine-tuning Hu et al. (2022), and single-stage versus multi-stage training methods. For details, see the Appendix F.
4.2 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a).
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities. Appendix G shows the details of the data division and usage.
4.3 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs include CuteGPT-13B [^9] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al. (2023), ChatGLM-6B [^10] https://huggingface.co/THUDM/chatglm3-6b Du et al. (2022), and Qwen-7B-Chat [^11] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al. (2023). These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
4.4 Metrics
The effectiveness of supplemental training will be assessed using the following metrics in the main experiments:
ACC: it represents the accuracy of the answer of LLM.
W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
To analyze the potential performance degradation caused by catastrophic forgetting, we also use the following metrics:
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{k}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of taks $i$ after applying CEM on task $k$.
5 Experiments
We construct supplemental training sets and control groups for experiments as outlined in Section 3.3. To differentiate models trained on various supplemental sets, we employ the following naming conventions:
INIT represents the model pre-finetuned solely with the Normative Corpus.
CEM-P: ‘P’ for Plain, indicates the model trained using the supplemental training set constructed by the Plain Strategy.
[…]
Suffixes ‘Wiki’, ‘Bing’, and ‘Mix’ indicate the data sources of the Supplemental Corpus, with ‘Mix’ signifying a combined set from Bing and Wikipedia. For instance, ‘EWiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
5.1 Main Results
We use the Plain Strategy to construct supplemental training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM. The main results are shown in Figure 3. Our observations are summarized as follows:
Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
Firstly, after supplemental training, a marked improvement in accuracy is noted across all models for both tasks. This suggests that the CEM method effectively addresses the model’s knowledge deficiencies, despite the limited training set and readily accessible supplemental data. For example, the ChatGLM-6B exhibits an increase of up to 16.08% in accuracy on the CMMLU task.
Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs. For example, on the Xiezhi task, ChatGLM-6B, with the fewest parameters, improves by up to 13.17%. In contrast, CuteGPT-13B, with the highest initial accuracy and most parameters, shows a smaller improvement of up to 3.63%.
Construction Strategy Analysis. To investigate the effectiveness of each component in the supplemental training set, we conducted experiments on CuteGPT-13B with various dataset construction strategies. We set the volume of both Extractive Corpus and Review Corpus to 1,000. The results are shown in Figure 4. Below is a summary of our observations:
[…]
Firstly, Extractive and Review Corpus significantly enhance LLM performance. CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics. Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric.
The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
Secondly, a Supplemental Corpus from multiple sources does not invariably enhance the CEM method more than a single source. Results in Figure 4 indicate no substantial difference in CEM effectiveness, whether using Wikipedia, Bing, or their combination. Notably, RMix records lower accuracy compared to RWiki and RBing, despite a larger dataset. This may arise from data redundancies. Moreover, an excessive supplemental corpus can reduce the proportion of CIT data, undermining its role in mitigating catastrophic forgetting, improving knowledge extraction, and increasing the risk of overfitting.
Thirdly, proportion factor $\boldsymbol{\alpha}$ has limited impact on the effectiveness of Review Corpus. Table 9 shows that R0, R0.5, and R1 have similar effects on LLM performance. This suggests that reviewing the knowledge in any format consolidates it. Therefore, unless specified, R will refer to the R0.5 in the following.
5.2 Multiple Iterations Analysis
The CEM method consistently yields significant improvements across multiple rounds, underscoring its iterative potential. We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, employing the same training strategies and data sources as in the first round. Independent CIT data of equivalent volume are used. Results in Table 10 indicate that this round significantly enhances the LLM’s accuracy across all dataset construction strategies. Notably, RWiki shows a 4.75% increase in accuracy over the previous round and a 17.00% improvement over INIT.
Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies. For example, the R2W metric of RWiki decreases by 7.02% compared to PWiki and by 5.90% compared to EWiki. We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplemental training.
Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM. The left part represents results without replay, while the right part shows results with a random replay strategy. ER indicates the enhancement rate on the Xiezhi task, and AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
Catastrophic Forgetting Analysis
To analyze catastrophic forgetting, we test CuteGPT-13B on Xiezhi, HotpotQA, and GSM8K to evaluate domain knowledge, reasoning, and mathematical skills. We use the CEM method to improve performance on Xiezhi and monitor any declines in the other tasks. Initially, we sample 2,000 instances from each dataset to establish a baseline. To mitigate forgetting, we employ random replay by replacing the General Corpus with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
Figure 5 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend. The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70.0% in most cases. However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by the random replay strategy), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 54.4% in round three without replay. CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect. Combining CEM with other advanced replay-based methods warrants further investigation.
6 Discussion
Why can the Extractive Corpus be effective? LLMs using the Extractive Corpus show higher accuracy and W2R metrics. Unlike the Review Corpus, where the role in performance improvement is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, while the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions act as a structured induction of the supplemental material, where the Extractive Corpus explicitly links complex information to the resulting induction (i.e., the question and options). Questions and options can serve as cues and guides for the model to gather relevant information from the corpus, helping it consider how this knowledge is accessed through questions. This process enhances the model’s ability to extract useful information from the corpus while filtering out redundant and irrelevant details.
7 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement. To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations.
8 Limitations
This paper focuses on the model’s performance in question-answering tasks. It remains uncertain whether CEM can consistently enhance the model’s capabilities and performance in non-QA tasks.
Future research will focus on investigating additional rounds of supplemental training to further enhance the LLM and explore the integration of other strategies to mitigate catastrophic forgetting. Additionally, it will examine whether repeated supplemental learning on already mastered corpora could have adverse effects.
Furthermore, we will experiment with diverse data sources, including synthetic corpora, and optimize strategies for collecting supplementary corpora to endow the LLM with advanced continual iterative evolutionary capabilities.
Appendix A Knowledge Point Extraction
Table 1 shows the prompt used for extracting knowledge points from questions using Qwen2-7B-Instruct. In the experiment, the variable $x$ in the prompt is set to $5$.
An Example of Knowledge Point Extraction
Instruction:
(task)Please analyze the core knowledge points examined by the following question.
(question)### Question:
What is the main component of the cell membrane?
A. Phospholipids B. Carbohydrates C. Proteins D. Nucleus
(answer)### Answer
The answer is A. Phospholipids. Phospholipids are the main component of the cell membrane.
(requirement)### Requirements:
- Prioritize identifying directly relevant named entities from the question and options.
- Knowledge points should be closely related to the question and options, aiding in eliminating incorrect options and selecting the correct one.
- Knowledge points should be specific, avoiding overly broad, common, or indistinct concepts, e.g., "Prime Minister", "Archaeology".
- Return the knowledge points in a list format, e.g., ["English Civil War", "Glorious Revolution"].
- The number of returned knowledge points should not exceed {x}.
Output:
(answer)["Cell Membrane Structure", "Phospholipid Bilayer"]
Table 1: An example illustrating the extraction of knowledge points from questions.
Appendix B Effectiveness Check of Supplemental Corpus
Table 5 shows the prompt used for entailment checking of the Supplemental Corpus with Qwen2-72B-Instruct. The results are presented in Table 2, where the proportion of useful information from both data sources consistently exceeds 65%.
Data Source Data Volume Effectiveness
Bing 8161 0.6718
Wiki 7694 0.6508
Table 2: The results of verifying whether the Supplemental Corpus contains information that helps the LLM in correcting errors and answering questions accurately. ‘Effectiveness’ denotes the proportion of useful information.
Appendix C Process of Collecting Corpora from Multiple Sources
The parameter configuration for the experiment is as follows: the value of $k$ is set to $4$ for top-k retrieval. Indexing is implemented using FAISS’s IndexFlatIP. The threshold for inner product similarity is set at $0.80$. If no entries meet this threshold, the entry with the highest similarity exceeding $0.70$ is selected; otherwise, no result is returned.
Fig.6 and Fig.7 show the process of obtaining supplemental data from Bing and Wikipedia respectively.
Figure 6: The process of collecting supplemental corpus from Bing.
Figure 7: The process of collecting supplemental corpus from Wikipedia.
Appendix D Format of CIT Data
Table 3 and Table 4 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
An Example of Normative Corpus
Instruction:
(task)Please select the correct answer for the following single choice questions:
(question)### Question:
In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
Output:
(answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 3: An example illustrating the format of the Normative Corpus.
An Example of Extractive Corpus
Instruction:
(task)Please select the correct answer for the following single choice questions based on the supplementary materials provided:
(question)### Question:
In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
(supplement)### Supplementary Materials:
The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history.
Output:
(answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 4: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix E Entailment Checking of Extractive Corpus
Table 5 shows the prompt used for entailment checking of the Extractive Corpus with Qwen2-72B-Instruct.
An Example of Entailment Checking
Instruction:
(task)Can the correct answer to the given multiple-choice question be derived solely from the knowledge in the following material? Please respond with ’Yes’ or ’No’.
(question)### Question:
What is the main component of the cell membrane?
A. Phospholipids B. Carbohydrates C. Proteins D. Nucleus
(answer)### Answer
The answer is A. Phospholipids. Phospholipids are the main component of the cell membrane.
(supplement)### Materials:
The phospholipid bilayer is a thin polar membrane made of two layers of lipid molecules. These membranes are flat sheets that form a continuous barrier around all cells. The cell membranes of almost all organisms and many viruses are made of a lipid bilayer, as are the nuclear membrane surrounding the cell nucleus, and membranes of the membrane-bound organelles in the cell…
Output:
(answer) Yes.
Table 5: An example illustrating the entailment checking of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix F Exploring Fine-Tuning Strategies
We experiment with fine-tuning strategies on Qwen-7B-Chat, employing full fine-tuning and LoRA fine-tuning for parameter updates. For dataset usage, we apply two approaches: single-stage parallel training on combined CPT and CIT data (3000 and 2000 samples, respectively), and multi-stage sequential training, where the model is first fine-tuned on CPT data, followed by further fine-tuning on CIT data.
Table 6 shows that parallel full fine-tuning yields the best results. Therefore, in the experiments, except for ChatGLM-6B, which employs Prompt-tuning Liu et al. (2022), all other models use parallel full fine-tuning.
Pre-Finetuning Parallel Sequential
Full 0.4162 0.4854 0.3154
LoRA - 0.4573 0.4271
Table 6: Results of verifying the impact of CPT and CIT datasets on model performance using different parameter update strategies and training stages.
Appendix G Data Division
Table 7 shows the the data division of datasets. For Xiezhi and CMMLU, the multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a normative format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing Supplemental Training Random Replay
Xiezhi 27,265 10,000 10,000 7,265 -
CMMLU 11,917 4,000 5,000 2,917 -
HotpotQA 7,000 2,000 5,000 - 500
GSM8K 7,000 2,000 5,000 - 500
Table 7: The data division and usage of Xiezhi, CMMLU, HotpotQA, and GSM8K. For HotpotQA and GSM8K, the data used for Random Replay is included in pre-finetuning.
Appendix H Experimental Result
Table 8 presents the comprehensive experimental results for various LLMs employing the CEM method on the Xiezhi and CMMLU tasks, with the supplemental training dataset constructed using the Plain Strategy.
Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
Table 10 presents the comprehensive experimental results of CuteGPT-13B on the Xiezhi task following two rounds of CEM, with the second round employing the same training data construction strategy as the first.
Table 11 presents the comprehensive results of catastrophic forgetting experiments for CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
Task Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Xiezhi Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
CMMLU Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
Table 8: Experimental results of various LLMs using the CEM method on the Xiezhi and CMMLU tasks
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0 CEM-R0.5 CEM-R1 CEM-R’
Wiki Acc 42.93 46.56 52.03 50.69 53.40 55.18 54.67 49.44
W2R - 22.20 24.12 23.42 24.82 24.93 24.61 22.90
R2W - 18.57 15.02 15.65 14.35 12.68 12.87 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.90 53.86 54.53 46.58
W2R - 21.91 23.15 21.37 24.99 24.25 25.18 22.28
R2W - 18.93 17.61 19.18 14.02 13.32 13.58 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.37 53.42 54.43 53.07
W2R - 21.99 25.12 23.67 25.00 24.04 25.06 24.81
R2W - 18.31 14.16 15.81 14.56 13.55 13.46 14.67
Table 9: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
Metrics Wiki Bing
CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 10: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task, where "v.s. INIT" indicates W2R and R2W calculated based on the initial evaluation results, and "v.s. Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
Round Strategy Xiezhi GSM8K HotpotQA ER AFR
INIT 40.84 13.20 80.60 1.00 1.00
Round 1 CEM-P 42.02 6.64 51.60 1.03 0.730
CEM-E 42.76 8.36 55.52 1.05 0.681
CEM-R 51.49 7.54 55.62 1.26 0.673
Round 2 CEM-P 48.68 9.94 57.84 1.19 0.723
CEM-E 49.47 9.76 60.24 1.21 0.746
CEM-R 54.23 6.88 56.68 1.33 0.703
Round 3 CEM-P 50.58 10.02 60.96 1.24 0.757
CEM-E 52.69 9.46 62.98 1.29 0.772
CEM-R 58.91 7.48 59.86 1.44 0.678
Table 11: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi, GSM8K, and HotpotQA tasks.